25. Suppose that in Bangladesh, three accidents occur in different factories per year on average. a. What is the probability of no accidents during a one-year period? 

b. What is the probability of at least two accidents during a one-year period? c. What is the probability of no accidents during the next six months?




1
Expert's answer
2021-05-24T00:36:01-0400

Poisson Distribution formula

 P(x; μ) = (e) (μx) / x!

where

x=Poisson random variable (x)

μ=Average rate of success=3 (given in question)


 a. What is the probability of no accidents during a one-year period? 

sol.

Poisson random variable (x)=0

Average rate of success=3

 P(x; μ) = (e) (μx) / x!

 P(0; 3) = (e-3) (30) / 0!

= 0.04978706836



b. What is the probability of at least two accidents during a one-year period?

sol.

Poisson random variable (x)=2 at least

Average rate of success=3

in the case of at least variable , we find this as

1-[   P(0; 3) +    P(1; 3) ]

1-[  (e-3) (30) / 0! } +   (e-3) (31) / 1! } ]

=0.80085



c. What is the probability of no accidents during the next six months?

sol.

Poisson random variable (x)=0

Average rate of success=3/2 (accidents in six months)

 P(x; μ) = (e) (μx) / x!

=0.22313